OpenStudy (anonymous):

Jada is trying to determine the number of t-shirts that were in inventory before she started her shift. During her shift, she sold thirteen t-shirts, which left twenty-two t-shirts in inventory. -Define a variable. -Write an equation Jada can use to determine the number of t-shirts that were in inventory before she started her shift. - Solve the equation showing your work.

OpenStudy (anonymous):

someone plz help

OpenStudy (anonymous):

So let's say the number of shirts are x. if she sold 13 shirts and still has 22 that means that x-13 = 22; x = 22 + 13; x = 35.

OpenStudy (anonymous):

thank you @JonyS